A perpetuity paying 1.000$ at the end of each month is replaced with an annuity paying X$ each month for 10 years. Calculate X if j4 = 7%


If the annuity paid 2.500$ at the end of each month, how long would the annuity last and what would be the size of the final, smaller payment?


1
Expert's answer
2021-11-08T06:14:38-0500

Solution:

j4=7%

i = 7/4 % = 1.75% p.a = 0.0175

PV of perpetuity "=\dfrac{\text{Amount of continuous cash payment}}{\text{Interest rate or yield}}"

"=\dfrac{\$ 1}{0.0175}=\$ 57\dfrac17"

PV of annuity "=P[\dfrac{1-(1+r)^{-n}}{r}]"

Given, "P=\$X, r=0.0175,n=10"

So, PV of annuity "=X[\dfrac{1-(1+0.0175)^{-10}}{0.0175}]=9.10122X"

Now, "9.10122X=57\dfrac 17"

"\Rightarrow X=\$6.278"


Now, "P=\$2.5,n=?"

PV of annuity "=P[\dfrac{1-(1+r)^{-n}}{r}]"

"PV=\$ 57\dfrac17"

"57\dfrac17=2.5[\dfrac{1-(1+0.0175)^{-n}}{0.0175}]
\\\Rightarrow 0.4=1-(1+0.0175)^{-n}
\\\Rightarrow (1.0175)^{-n}=0.6
\\\Rightarrow -n\log 1.0175=\log 0.6
\\\Rightarrow n=29.44\approx 30"


Now, for n=30,

"PV=2.5[\dfrac{1-(1+0.0175)^{-30}}{0.0175}]=57.964"

Final payment will be "=57.964-57\dfrac17=\$0.823"